Common collector amplifier has an input resistance of Rin=10kΩ and load 

resistance of RL=100Ω. Power gain for this amplifier is 100. What is the voltage gain 

of the amplifier?


1
Expert's answer
2021-06-23T13:09:05-0400

Gives

"R_{in}=10K\Omega"

"R_L=100\Omega"

"A_p=100"

"A_p=A_iA_v"

"A_v=\frac{A_p}{A_i}"

Put value

"A_v=\frac{100}{A_i}"

"A_i=\frac{R_L}{R_i}"


"A_v=\frac{100}{A_i}=\frac{100}{\frac{R_L}{R_i}}=\frac{100\times10000}{100}=10^4"

"A_v=10^4"